# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

### v1

@@ Abstract
In order to compute the Fourier transform of a function $f$ on the real line numerically, one samples $f$ on a grid and then takes the discrete Fourier transform. We derive exact error estimates for this procedure in terms of the decay and smoothness of $f$. The analysis provides a new recipe of how to relate the number of samples, the sampling interval, and the grid size.
@@ 1 Introduction
$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}x\xi}\mathrm{d}x$ (1)
@@
    Sample $f$ on an interval of length $p$ at $n$ equispaced points $hj$, for $j\in\mathbb{Z}$, $-\frac{n}{2}<j\leq\frac{n}{2}$, with step size $h$.
@@
    It is then taken for granted that the scaled discrete Fourier transform of length $n$ provides an approximation for the values of the continuous Fourier transform of $f$ at $k/p$, i.e.,
    $\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\,,\qquad k\in\mathbb{Z},\quad-\frac{n}{2}<k\leq\frac{n}{2}\,.$ (2)
This practice is usually unquestioned and works like a charm. It is motivated by following approximations with quadrature and truncation,
$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}\xi x}\mathrm{d}x\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\,.$ (3)
For $\xi=\frac{k}{p}=\frac{k}{hn}$, the right-hand side is the scaled discrete Fourier transform (2). Yet one must ask what is the precise relation between the computed values $h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}$ and the samples $\hat{f}(\frac{k}{p})$ of the actual Fourier transform. Ideally one can provide reasonable error estimates for this approximation, and we formulate two questions:
@@
    The numerical procedure depends on the number of samples $n$, the step size $h$, and the length $p$ of the sampled interval via the identity $p=hn$. How should $h,n$ and $p$ be chosen in dependence of the function class of $f$?
@@
    What can be said about the asymptotic decay of the deviation in (2) when $n,p\rightarrow\infty$ and $h\to 0$?
State-of-the-art. To the best of our knowledge, there are surprisingly few investigations and even fewer quantitative error estimates in the literature.
The “Owner’s Manual for the Discrete Fourier Transform” [4, Section 6] contains a substantial chapter on the error of the discrete Fourier transform. There the error is quantified for rather special function classes, e.g., functions with compact support or functions on $\mathbb{R}$ with exponential decay with some type of finite order of smoothness conditions [4, Theorem 6.6] that are rarely met, see also our Section 3.5.
[…]
@@
Since the right-hand side of (2) approximates the integral $\int_{-p/2}^{p/2}f(x)e^{-2\pi ikx/n}\mathrm{d}x$ by means of a Riemann sum or the trapezoidal rule, one should mention the recent error analyses [7, 11, 23, 25, 24] for the trapezoidal rule as relevant in this context. However, these error estimates require exponential decay and analyticity and do not even mention the Fourier transform.
@@
In this paper we derive estimates for the error between the Fourier transform on the real line and the discrete Fourier transform for a large class of functions with polynomial or sub-exponential decay and finite smoothness. Our results confirm that, even under such very mild assumptions on the decay and smoothness, the folklore procedure works well and is successful. Our analysis also provides the new insight that the optimal spacing should depend on the decay of $f$ and $\hat{f}$. In view of (Q1), we identify an optimal relation between the step size $h$, the number of samples $n$, and the length of the sampling interval $p$ (the step size $\frac{1}{p}$ in frequency). To answer (Q2), we derive precise error estimates with explicit constants for function classes that are significantly larger than exponentially decaying, analytic functions. These error estimates are confirmed by numerical simulations utilizing the FFT which yield the precise asymptotics predicted by our theoretical estimates.
Our contribution. We proceed with a formal presentation of the numerical procedure and an explicit exposition of our results.
@@
$[n]:=\{j\in\mathbb{Z}:-\tfrac{n}{2}<j\leq\tfrac{n}{2}\}\,.$
@@
$\operatorname{\mathcal{F}}\!y=\bigg{(}\frac{1}{\sqrt{n}}\sum_{j\in[n]}y_{j}e^{-2\pi\mathrm{i}\frac{kj}{n}}\bigg{)}_{k\in[n]}\,.$ (4)
@@
Throughout the text we use the relation $p=hn$ between the interval length $p$, the step size $h$ and the number of samples $n$.
Our aim is to bound the approximation error
$E^{[n]}_{h}(f):=\biggl{(}\frac{1}{p}\sum_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}^{2}\bigg{)}^{1/2}=\big{\|}\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big{\|}\,.$ (5)
The normalization of $f_{h,n}$ and the factor $\frac{1}{p}$ in the definition of the error may seem artificial, but they are justified and natural in view of the time-frequency symmetry of the error
$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
See Appendix A for the proof. The symmetry between $f$ and $\hat{f}$ will also appear in the error estimates.
The following error estimate is a simple consequence of our main result for polynomial decay.
@@
Let $a,b>1$. If $f$ and $\hat{f}$ are continuous and satisfy polynomial decay of the form [^1] We write $\lesssim$ if the left-hand-side is bounded by a constant times the right-hand-side. If $\lesssim$ and $\gtrsim$ both hold, then we write $\asymp$. The dependency of the constants on other parameters shall be clarified or is clear from the context. $|f(x)|\lesssim(1+|x|)^{-a}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$, then for all $\alpha<a-\frac{1}{2}$, $\beta<b-\frac{1}{2}$
$E^{[n]}_{h}(f)\lesssim p^{-\alpha}\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,,$
where the constants in $\lesssim$ are independent of $h$, $n$, $p$.
The error bound is balanced if $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$, and for this choice $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$. This is the optimal choice for the relation between $h$, $p$, and $n$, see Corollary 3.9. Our numerical simulations in Section 2 using the FFT suggest that the bounds in Theorem 1.1 are sharp.
For the approximation procedure to make sense, we need the Fourier transform to be defined pointwise and continuous, so that $f\in L^{1}(\mathbb{R})$ is a reasonable minimal assumption. One may argue that we also make certain assumptions on $\hat{f}$, which is the very object we want to compute. If one does not like the polynomial decay condition of $\hat{f}$ in Theorem 1.1, one can replace it by a stronger condition on smoothness of $f$. Precisely, if the derivatives $f^{(l)}$ are in $L^{1}(\mathbb{R})$ for $l=0,\dots,b$ with $b\in\mathbb{N}$, then it follows that $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$.
Sub-exponential decay in time and in frequency occurs importantly in the theory of test functions and distributions (under the name of Gelfand-Shilov spaces or ultra test functions) [3]. In that case, we offer the following version.
@@
Let $r,s>0$ and $0<\alpha,\beta\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy the decay conditions $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim e^{-s|x|^{\beta}}$, then for all $r^{\prime}<r$,s<$s^{\prime}<s$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}+e^{-s^{\prime}(2h)^{-\beta}}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|e^{s|\xi|^{\beta}}\,.$
@@
We also provide bounds for mixed decay.

### v2

@@ Abstract
In order to compute the Fourier transform of a function $f$ on the real line numerically, one samples $f$ on a grid and then takes the discrete Fourier transform. We derive exact error estimates for this procedure in terms of the decay and smoothness of $f$. The analysis provides an asymptotically optimal recipe of how to relate the number of samples, the sampling interval, and the grid size.
@@ 1 Introduction
(1) $\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}x\xi}\mathrm{d}x$
@@ 1.1 Questions
    Sample $f$ on an interval $(-\frac{p}{2},\frac{p}{2}]$ of length $p$ at $n$ equispaced points $hj$, for $j\in\mathbb{Z}$, $\frac{n}{2}<j\leq\frac{n}{2}$, with step size $h$ and $p=hn$.
@@
    Approximate the Fourier transform of $f$ at $\tfrac{k}{p}$ by
    (2) $\hat{f}\big(\tfrac{k}{p}\big)\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\,,\qquad k\in\mathbb{Z},\quad-\frac{n}{2}<k\leq\frac{n}{2}\,.$
The right-hand side is the discrete Fourier transform of the sampled vector $\big(f(hj)\big)_{j}$ of size $n$. By choosing $n$ to be a power of $2$ or by zero-padding, the discrete Fourier transform can be computed with the FFT algorithm in $\mathcal{O}(n\log n)$ operations. Since the Fourier transform is a fundamental computational step in a huge number of applications in signal processing, acoustics, medical imaging, the numerical solution of partial differential equations, in quantum mechanics, etc., the availability of a fast algorithm is central in scientific computing. Therefore much effort has gone into a speed up and efficient implementations of this algorithm [25, Chapter 12]. The FFT is hailed as “the most important numerical algorithm of our life-time” [29], is included in the ten most important numerical algorithms of the 20th century [9], and remains important in the age of data science and artificial intelligence.
Grounded on the overwhelming success of the FFT in approximating the Fourier transform, the approximation (2) is usually taken for granted. The FFT is used as a black box, it works superbly, and it always seems to work.
Qualitatively, the approximation (2) is easy to understand. It is motivated by discretizing the integral and then truncating the infinite series as follows:
(3) $\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}\xi x}\mathrm{d}x\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\,.$
For $\xi=\frac{k}{hn}=\frac{k}{p}$ one then obtains (2).
Although this approximation is immensely successful, one must ask what is the precise relation between the computed values $h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}$ and the samples $\hat{f}(\frac{k}{p})$ of the actual Fourier transform.
Ideally one can provide reasonable error estimates for this approximation, and we formulate two questions:
@@
    The numerical procedure depends on the number of samples $n$, the step size $h$, and the length $p$ of the sampled interval where $p=hn$. How should $h,n$, and $p$ be chosen dependent on the function class of $f$?
@@
    What can be said about error rates and the asymptotic decay of the deviation in (2) when $n\rightarrow\infty$, $p\rightarrow\infty$, and $h\to 0$?
We will answer these questions for several classes of functions that are described by (i) their decay property and (ii) by their smoothness or by the decay of their Fourier transform. The decay of $f$ determines the truncation error in (3), and the decay of $\hat{f}$ determines the discretization error in (3).
1.2 State-of-the-art
The case of functions with compact support is related to the case of approximating Fourier coefficients of a periodic function and is fully understood. Roughly, if $f$ has compact support and is $m$-times differentiable, then the pointwise error scales like $n^{-m}$ in terms of length of the FFT, or like $h^{m}$ in terms of the grid size. This is made precise by Epstein in [11], whose title we borrowed, for the approximation of Fourier coefficients of periodic functions, and by Briggs and Henson in their “Owner’s Manual for the Discrete Fourier Transform” [5, Section 6] for the Fourier transform of compactly supported functions on $\mathbb{R}$. For $\operatorname*{supp}f\subseteq[-\frac{p}{2},\frac{p}{2}]$ and $h=\frac{p}{n}$, these results are optimal and come with explicit constants.
The article [2] computes an exact formula for the relative pointwise error for the Fourier transform of a compactly supported $B$-spline of order $k$ (called “canonicalk-k functions”) and shows that in the limit $n\to\infty$ the relative error is the same for $k+1$-times differentiable functions with compact support, however, no error estimates are given.
To the best of our knowledge, there are surprisingly few investigations and even fewer quantitative error estimates in the literature that go beyond compactly supported functions. For functions with unbounded support the important cases of exponential decay and analyticity of $f$ are covered by the work of Briggs and Henson [5, Chapter 6] and of Stenger [28, Chapter 3.3].
The case of general input in [5, Thm. 6.6] assumes that $f$ satisfies $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}<\infty$ and is $m$-times differentiable satisfying some stringent boundary conditions at $\pm p/2$. Then for some constants $c_{1},c_{2}$ the pointwise error is
(4) $\Big|\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq c_{1}e^{-rp/2}e^{-rp/n}+c_{2}h^{m},\qquad-\frac{n}{2}<k\leq\frac{n}{2}\,.$
The decay of order $h^{m}$ is as expected, and the implicit assumption of bounded variation [5] even leads to $h^{m+1}$. The problem, however, is in the boundary conditions. In the generic case, when the values of $f$ at the endpoints of the sampled interval differ, i.e., $f(-p/2)\neq f(p/2)$, then the predicted error is only of the order $h$. For instance, for the exponential function $f(x)=e^{-2\pi|x|}$ [5] obtains the correct error estimate with $h^{2}$, but for the shifted function $e^{-2\pi|x-1|}$ the predicted error is only $h^{1}$, which does not match the numerical observations. This phenomenon seems to be a limitation of the proof technique in [5]. It is one of our contributions to obtain the correct error estimate without any boundary conditions. For more discussion see our Section 4.7.
In the context of numerical sinc methods Stenger [28, Chapter 3.3] uses the cardinal series to approximate an exponentially decaying function $f$ whose Fourier transform also decays exponentially (to be precise, he uses a slightly stronger condition formulated with a Hardy space) and shows that the pointwise error satisfies
$\Big|\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq Ce^{-sn^{1/2}}\,$,
with suitable constants $C,s>0$. In addition, [28, Chapter 3.3] contains error estimates for the $L^{p}$-error for interpolated versions of the Fourier transform.
@@ 1.2 State-of-the-art
Since the right-hand side of (2) approximates the integral $\int_{-p/2}^{p/2}f(x)e^{-2\pi\mathrm{i}xk/p}\mathrm{d}x$ by means of a Riemann sum, one should mention the recent error analyses [10, 14, 30, 32, 31] for the trapezoidal rule as relevant in this context although they do not mention the Fourier transform. These error estimates assume exponential decay and analyticity of $f$. By contrast, the recent results for the trapezoidal rule in [21, Prop. 4.2, Thm 4.5] do not require analyticity, but use Gaussian decay and finite Sobolev smoothness of order $m$. For $n$ samples of $f$ and the choice $p=2\sqrt{\frac{2}{1-\epsilon}m\log n}$ the error is of the order $n^{-m}(\log n)^{m/2+1/4}$.
In [8] the trapezoidal rule is used for error estimates of approximations of the Fourier transform of analytic functions via doubly exponential transforms. These are rather different from the usual approximation (2), but yield almost exponential convergence.
An interesting variation was studied by Auslander and Grünbaum [1]. They consider functions $f\in L^{2}(\mathbb{R})$ and approximate averaged values of $\hat{f}$ from local averages of $f$ via the discrete Fourier transform. For Gaussian smoothing they derive an explicit formula for the sharp constant for the resulting error estimate and then plot the results of numerical simulations for FFTs up to $n=100$ samples. Asymptotics for large $n$ is not treated.
2 Our contributions
In this paper we derive estimates for the error between the Fourier transform on the real line and the discrete Fourier transform for a significantly broader class of functions than previously considered in the literature.
@@ 2 Our contributions
Specifically, we study functions that decay only polynomially in time and in frequency. Our results confirm in a quantitative fashion that, even under mild assumptions on the decay and smoothness (decay in frequency), the standard approximation procedure with the FFT works well and is successful. Our analysis also provides the new insight how the optimal spacing should depend on the decay of $f$ and $\hat{f}$. As an answer to Question (Q1), we identify an optimal relation between the step size $h$, the number of samples $n$, and the length $p$ of the sampling interval. To answer Question (Q2), we derive precise error estimates with explicit constants for function classes that are significantly larger than exponentially decaying or analytic functions. The optimality of the error estimates is confirmed by numerical simulations with the FFT that yield the precise asymptotics predicted by theoretical results. We believe that these error estimates are best possible, and since all constants are explicit, these estimates cover not only the asymptotic regime, but also hold for small $n$.
We proceed with the detailed exposition of our results.
2.1 Error measures
@@
$n]:=\{j\in\mathbb{Z}:-\tfrac{n}{2}<j\leq\tfrac{n}{2}\}\,.$
@@
(5) $\operatorname{\mathcal{F}}\!y=\bigg(\frac{1}{\sqrt{n}}\sum_{j\in[n]}y_{j}e^{-2\pi\mathrm{i}\frac{kj}{n}}\bigg)_{k\in[n]}\,.$
@@
Throughout the text we use the relation $p=hn$ between the interval length $p$, the step size $h$, and the number of samples $n$.
We will deal mainly with the $\ell^{2}$-error averaged over the sampled interval. This approximation error is defined as
(6) $E^{[n]}_{h}(f):=\biggl(\frac{1}{p}\sum_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|^{2}\bigg)^{1/2}=\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big\|\,.$
The use of the $\ell^{2}$-error is natural in our context, since the discrete Fourier transform is unitary on $\mathbb{C}^{n}$ with respect to the $\ell^{2}$-norm. The normalization of $f_{h,n}$ and the factor $\frac{1}{p}$ are motivated by the remarkable time-frequency symmetry of the error
[…]
see Lemma 4.1. This symmetry aligns well with the mathematical framework of time-frequency analysis in [17] in which both time and frequency are treated as equally important dimensions for understanding signals. The symmetry between time and frequency also manifests in our error estimates, and we believe that the unified time-frequency perspective is a key reason why our estimates are sharp.
2.2 Polynomial decay and finite smoothness
Our main theorem will be proved for general decay conditions. We first discuss the special case of functions of polynomial decay and finite smoothness. As remarked in  [5], finite smoothness is often replaced by polynomial decay of the Fourier transform, which we will do here. The first error estimate is a simple consequence of our main result (Theorems 4.7 and 4.6) and is the most relevant special case.
@@ 2.2 Polynomial decay and finite smoothness
Let $a,b>1$. If $f$ and $\hat{f}$ are continuous and satisfy polynomial decay of the form $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$, then, for all $\alpha<a-\frac{1}{2}$ and $\beta<b-\frac{1}{2}$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
(7) $E^{[n]}_{h}(f)\leq c\,(p^{-\alpha}+h^{\beta})\,.$
This statement can be read on several levels.
  1. (i)
    It confirms the users’ experience that the discrete Fourier transform indeed approximates the continuous Fourier transform.
  2. (ii)
    Theorem 2.1 offers guarantees of how well the FFT approximates the Fourier transform by quantifying the approximation error in terms of the decay of $f$ and $\hat{f}$.
  3. (iii)
    The theorem provides a hint on how to improve the accuracy of the numerical computations. This is a new insight with possibly practical consequences.
    If no a priori knowledge of $f$ is given, then the $n$ samples of $f$ are usually chosen on an interval of length $p=\sqrt{n}$ with grid size $h=1/{\sqrt{n}}$. In this case the error decays like $E^{[n]}_{h}(f)=\mathcal{O}(n^{-\min(\alpha,\beta)/2})$. However, with some knowledge about these decay rates, one may balance the two terms in (7) so that both terms contribute equally***We write $\lesssim$ if the left-hand-side is bounded by a constant times the right-hand-side. If $\lesssim$ and $\gtrsim$ both hold, then we write $\asymp$. The dependency of the constants on other parameters shall be clarified or is clear from the context., i.e., $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$. For this choice the error is
    (8) $E^{[n]}_{h}(f)\leq cn^{-\frac{\alpha\beta}{\alpha+\beta}}$
    This observation will become relevant when the decay rates $\alpha,\beta$ are known and the difference $|\alpha-\beta|$ is large. The choice $p^{-\alpha}\asymp h^{\beta}$ then leads to a significantly more accurate numerical approximation of the Fourier transform using the same number of samples. Conversely, a given target level of accuracy can be achieved with significantly fewer samples. See Section 3.
    We also note that this choice leads to the asymptotically optimal relation between $h$, $p$, and $n$, see Corollary 4.11.
  4. (iv)
    In Corollary 4.12 we will determine the constants involved in the balancing $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$, and we will determine the constant $c$ in (8) explicitly in terms of $\alpha,\beta$ and a norm on $f$. We emphasize that the ultimate version of the error estimate is a non-asymptotic result that may yield relevant information about the error even for small $n$.
Our numerical simulations in Section 3 confirm that the bounds in Theorem 2.1 are sharp up to some infinitesimal margin.
Smoothness instead of Fourier decay
One may argue that we also make certain assumptions on $\hat{f}$, which is the very object we want to compute. The polynomial decay condition of $\hat{f}$ in Theorem 2.1 can be replaced by a stronger condition on smoothness of $f$.
Precisely, if the derivatives $f^{(l)}$ are in $L^{1}(\mathbb{R})$ for $l=0,\dots,b$ with $b\in\mathbb{N}$, then it follows that $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$. Using this observation, we obtain a slightly weaker, but equally useful version of Theorem 2.1.
Corollary 2.2.
Let $a,m>1$. Assume that $|f(x)|\lesssim(1+|x|)^{-a}$ and that $f$ has $m$ derivatives in $L^{1}(\mathbb{R})$. Then for $0<\alpha<a-1/2$ and $\beta<m-1/2$ there is a constant $c>0$ independent of $h$, $n$, $p$ such that
(9) $E^{[n]}_{h}(f)\leq c\,(p^{-\alpha}+h^{\beta})\,.$
2.3 Sub-exponential decay
Exponential and sub-exponential decay in time and in frequency occur importantly in the theory of test functions and distributions (under the name of Gelfand-Shilov spaces or ultra test functions) [3]. The strongest form is exponential decay for which error estimates have been treated before. We offer the following version of the error.
@@ 2.3 Sub-exponential decay
Let $r,s>0$ and $0<\alpha,\beta\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy the decay conditions $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|e^{s|\xi|^{\beta}}<\infty$, then for all $r^{\prime}<r$ and $s^{\prime}<s$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
$E^{[n]}_{h}(f)\leq c\,(e^{-r^{\prime}(\frac{p}{2})^{\alpha}}+e^{-s^{\prime}(2h)^{-\beta}})\,.$
@@
For the parameters $\alpha=\beta=1$, the assumption says that both $f$ and $\hat{f}$ decay exponentially. The choice $p=\sqrt{n}$ and $h=1/{\sqrt{n}}$ then yields the error estimate
$E^{[n]}_{h}(f)\leq c\,e^{-d\sqrt{n}}\,$,
with a constant $d$ depending on $r,s,\alpha,\beta$. Up to possibly different constants, this root-exponential convergence is compatible with Stenger’s result [28].
2.4 Mixed conditions
We finally provide bounds for mixed conditions when $f$ possesses exponential decay, but finite smoothness.